### v1

# Spatiotemporal representation learning for short and long medical image time series

Abstract
Analyzing temporal developments is crucial for the accurate prognosis of many medical conditions. Temporal changes that occur over short time scales are key to assessing the health of physiological functions, such as the cardiac cycle. Moreover, tracking longer term developments that occur over months or years in evolving processes, such as age-related macular degeneration (AMD), is essential for accurate prognosis. Despite the importance of both short and long term analysis to clinical decision making, they remain understudied in medical deep learning. State of the art methods for spatiotemporal representation learning, developed for short natural videos, prioritize the detection of temporal constants rather than temporal developments. Moreover, they do not account for varying time intervals between acquisitions, which are essential for contextualizing observed changes. To address these issues, we propose two approaches. First, we combine clip-level contrastive learning with a novel temporal embedding to adapt to irregular time series. Second, we propose masking and predicting latent frame representations of the temporal sequence. Our two approaches outperform all prior methods on temporally-dependent tasks including cardiac output estimation and three prognostic AMD tasks. Overall, this enables the automated analysis of temporal patterns which are typically overlooked in applications of deep learning to medicine.
1 Introduction
Analyzing temporal developments in medical images is crucial to the practice of medicine. Clinicians analyze temporal change over short time scales, such as arrhythmia in the cardiac cycle in MR videos [22, 3, 23] or prenatal movement in ultrasound [27], in order to analyze dynamic physiological processes and function. Clinicians also track the development of processes that evolve gradually, such as disease progression [21, 19, 18] or a patient’s response to treatment [5, 14], over long time frames. In this longitudinal data, images are acquired at sparse and typically irregular intervals that may be months or even years apart. Despite the importance of both temporal dynamics and historical observations to the present diagnosis and prognosis of patients, the development of representation learning methods capable of learning these trends in medical data has thus far been overlooked.
Current state of the art approaches for spatiotemporal learning [13, 28] were primarily developed on the Kinetics dataset [20], a human action recognition dataset which depicts only a single action or event per video. To predict this singular event, these methods prioritize the detection of video-level features that persist over time, and in doing so explicitly ignore changes between frames. As such, these methods are unsuitable for learning representations that capture variance over the timeline. Furthermore, these methods were designed for videos with a fixed framerate and are thus unable to account for the rate of change in longitudinal sequences with irregular sampling.
To address these issues, we contribute the following:
We identify a diverse set of medical tasks that require modeling temporal developments over both short and long time frames (see Figure 1). We model short term patterns in cardiac MR video, which depicts the cardiac cycle over 1.6 seconds in 50 frames, to make diagnostic and biometric predictions associated with heart function. We analyze long range temporal changes in longitudinal optical coherence tomography (OCT), predicting current and future developments in patients with age-related macular degeneration (AMD), from up to six years of scanning history acquired at irregular intervals.
We demonstrate that, while established strategies for representation learning in natural videos perform adequately on short cardiac videos, they are especially poor at modeling long range developments in longitudinal retinal images.
We then propose a simple clip-level contrastive learning strategy that uses fixed clips and a temporal embedding to encode varying time intervals, outperforming all natural video baselines.
Finally, we propose a novel method that combines the efficacy of clip-level contrastive learning with a frame-level latent feature prediction task. Our method, illustrated in Figure 2 also outperforms all natural video baselines in three prognostic tasks for late stage AMD.
2 Related Work
2.1 Self-Supervised Video Representation Learning
In the image domain, contrastive learning has emerged as a strong and robust method for learning global imaging features without labels [7, 15, 16, 8, 4]. Similarly, contrastive approaches have achieved the state-of-the-art on the standard benchmark dataset for natural videos, Kinetics [20]. By minimizing the feature distance between differing segments, or clips, sampled from different time points in the same video, contrastive approaches learn to extract the single human action recorded over the span of each Kinetics video [28, 13, 35]. Specifically, Qian et al. [28] adopt a probabilistic function to vary the time between the two sampled clips, Feichtenhofer et al. [13] sample up to four clips, and Wang et al. sample clips of varying lengths [35]. However, by learning features that persist across the entire video these methods ignore temporal patterns which are associated with many medical conditions, including those involved in longitudinal disease progression.
To address this limitation, some recent works modify the contrastive framework to instead maximize the distance between the sampled clips [10, 40], or by parameterizing the video with an evolving stochastic process [26, 38]. Yang et al. in Latent time navigation (LTN) [37] introduce an extra orthogonal basis that explicitly models the temporal shift between sampled clips of the same video. Generative approaches, such as VideoMAE [34] and v-JEPA [2], implicitly model temporal variations by masking and reconstructing a subset of frame-level features. It is unclear whether the ability of these methods to model temporal developments on natural videos translates to spatiotemporal medical data.
2.2 Representation Learning for Spatiotemporal Medical Data
In the medical field, spatiotemporal representation learning has many applications, such as organ segmentation [31], physiological development analysis [39], and disease progression prediction [9]. Some works leverage the assumption that certain disease progressions share similar trajectories in the latent space [25, 24]. In addition, [29] and [36] leverage redundancies between successive brain scans to improve segmentation accuracy. However, there is no unified approach that learns self-supervised representations of spatiotemporal data in the medical domain. In this work, we propose a method that combines the efficacy of contrastive approaches with a masked latent reconstruction loss to model trajectories of frame-level temporal changes that are crucial for downstream medical analysis.
3 Materials and Methods
3.1 Short and long time scale datasets
3.1.1 Cardiac MR video
To study temporal variation over short time frames we chose to use short-axis cardiac MR videos sourced from the UK Biobank (UKBB) [32] (see Figure 1). The dataset contains 46,137 videos where each captures at least one full contraction, from end diastole to end systole, in 50 frames taken over 1.6 seconds. After resampling the images to 1.8 mm in the sagittal and coronal dimensions, we used segmentation masks to crop a region of about the center of mass of the hearts.
3.1.2 Longitudinal retinal OCT
To study temporal change over long time frames using sparse, irregular data, we use an in-house retinal OCT dataset tracking patients with age-related macular degeneration (see Figure 1) collected by the Southampton Eye Unit in the scope of the PINNACLE study [33]. The dataset includes 48,825 OCT scans, monitoring 6,368 eyes in 3,498 patients. On average, each eye was scanned 7.7 times over 1.9 years, where eyes with AMD were observed for an average of 2.3 years.
3.2 Spatiotemporal encoder
As the basis of all baselines and new approaches, we use a spatiotemporal encoder, that works on series of images. The spatial encoder is a Vision Transformer (ViT) [11] that extracts a representation of each image independently using a learnable CLS token. These CLS tokens are then collectively input to a temporal transformer, which is able to model temporal change across the series timeline.
3.3 Clip-level contrastive learning
Established contrastive learning methods minimize the feature distance between multiple temporally distant clips of the same sequence [28, 13](see Figure 2 left). To model temporal variations rather than temporal constants, we construct a clip-level contrastive approach that augments only a single randomly sampled clip, . We then generate a pair of two augmented views of our clip, and , where the augmentations in each are applied consistently across timepoints (see Figure 2 middle). We pass these views through the spatiotemporal encoder to extract two clip-level contrastive tokens and , which are appended as learnable tokens to the input of the temporal encoder. Finally, we optimize the SimCLR NT-Xent loss over the samples in the batch using equation 1, where is the cosine similarity, is the temperature, is a non-linear projection head following the design of [7].
3.4 Time embedding
To capture the irregular time differences between scans, we propose time embeddings (TE) for transformers. For each clip, we calculate the relative time between each scan and the first scan. We then apply a learnable non-linear mapping based on Ho et al. [17] to map the relative times to the feature dimension of the model. Similar to a position embedding, we add these temporal embedding to the frame-level tokens. This enables the temporal encoder to estimate the rate of change in longitudinal sequences with irregular sampling (see Figure 2 middle).
3.5 Temporally-variant feature prediction
Finally, to explicitly model frame-level temporal variations, we propose a temporal feature prediction task inspired by Assran et al. [1] (see Figure 2 right). Specifically, given a sequence of frame-level tokens output by our spatial encoder, we replace a random subset with a learnable mask token using a masking ratio of 0.15. We then use a projection head on top of the temporal encoder output tokens to reconstruct the masked tokens. We then compute the cosine similarity loss using equation 2 between the reconstructed tokens and the original spatial tokens , where denotes the number of masked tokens.
The final loss, termed temporally-variant representation loss (TVRL), combines clip-level contrastive learning with frame-level feature prediction and is defined in equation 3, where is a weight to combine the two terms.
4 Experiments and Results
4.1 Experimental setup
4.1.1 Pretraining protocol
We evaluate two proposed approaches: the clip-level SimCLR strategy described in section 3.3, and our TVRL extension introduced in section 3.5 using . For both strategies, we then evaluate the effect of adding time embeddings, described in section 3.4. We compare these against the leading spatiotemporal representation learning strategies which includes CVRL [28], LTN [37], and VideoMAE [34]. All pretraining strategies use a spatial encoder parameterized by a ViT-S [11] with 384 hidden dimensions, six heads, and 12 attention layers, while the the temporal encoder uses the same configuration but with three layers of attention. In total the model has 27.7M parameters.
Before pretraining, we first partition each dataset into training, validation, and test sets, comprising 70%, 15%, and 15% of the data, respectively, ensuring that each patient appears in only one subset. Each model accepts eight images as input. We sample every other frame from the cardiac videos, and every successive scan in the retinal longitudes. We employ separate projection heads and for our contrastive and masked reconstruction losses, following the two-layer configuration in [7]. All models are pretrained for 200 epochs with a batch size of 256 using the AdamW optimizer, and a cosine annealing learning rate of with 20 epochs of linear warm-up.
4.1.2 Finetuning protocol
We follow the standard linear probing protocol [6] and finetune each pretrained model by freezing the network and appending a new trainable [CLS] token to the temporal transformer which serves as input to a single linear layer. We finetune each model with 1% and 100% of the data, and report the area under the receiver operating characteristic curve (AUC) for classification tasks and mean absolute error (MAE) for regression tasks, with mean and standard deviation over two random seeds. We finetune with AdamW for 100 epochs with a learning rate of and a batch size of 256 and report test performance of the best validation epoch. As baselines to the aforementioned strategies, we also evaluate a fully supervised spatiotemporal encoder trained end-to-end, in addition to a 3D ResNet50 [12] (31.6M parameters) pre-trained on the Kinetics-400 dataset [20].
4.2 Downstream evaluation
We first evaluate different pretraining strategies on a set of control tasks that are marginally associated with temporal change. In cardiac MR video, we predict coronary artery disease (CAD) and fibrillation from ICD codes recorded during hospital admissions of all subjects. From retinal OCT longitudinal data we predict the presence of late vs. early AMD given the present and historical set of scans of each patient.
We then evaluate the models on tasks that are strongly associated with temporal change. In the short time scale, we estimate left ventricular ejection fraction (LVEF) and total cardiac output, which measures the volume of blood pumped in one minute. Finally, we focus on three OCT-based prognostic tasks that require extrapolating long term trends from the patient’s history of disease progression. These involve predicting the development of late stage wet AMD (defined by choroidal neovascularisation), late stage dry AMD (defined by cRORA > 1000 [30]) and retinal scarring and fibrosis post treatment, all within the next 3 years.
4.3 Simple clip-level contrastive learning outperforms prior approaches
In Table 1 and Table 2, we find in all tasks that the time-invariant representations learned by CVRL [28] degrade performance compared to our simple time-dependent SimCLR [7] approach using a single clip. This was most pronounced on the most temporally dependent tasks such as prognosis of late dry AMD (76.6% vs. 68.4% AUC). While LTN [37] and VideoMAE [34] improve the modeling of temporal variations, both methods and baselines perform comparably to or worse than clip-level SimCLR across all tasks in both the short and long time scales. Results on cardiac tasks were comparable across most models.
4.4 Explicit temporal modeling improves long term prognostic tasks
The clip-level SimCLR approach with time embeddings (SimCLR + TE) outperforms all prior methods on the irregularly sampled retinal OCT longitudes. It improved the prognosis of late dry AMD from 76.6% to 81.6% AUC and scarring from 72.1% to 79.2% AUC, as shown in Table 1. It also performed best in estimating the cardiac output, achieving 1.63 MAE as shown in Table 2. Finally, TVRL further boosts the performance on prognosis for wet AMD to 68.8% AUC and scarring to 80.5% AUC with time embedding, while reaching comparable performance as clip-level SimCLR in predicting Late vs. Early AMD.
5 Discussion and Conclusion
In this paper, we identified two common variants of spatiotemporal data used in clinical practice. We found that while state-of-the-art spatiotemporal learning methods developed for natural videos perform comparably in short term analysis on cardiac MR video, they failed to track long term developments and disease trajectories in longitudinal series, leading to poor prognostic utility for AMD. To address this, we propose a simple clip-level contrastive learning strategy that leverages time embeddings in irregular and variable length time series, and a new temporally-variant approach that explicitly models frame-level variation. Models pretrained with our strategies had improved assessment of cardiac output, and substantially improved prognosis for three variants of late stage AMD.
One limitation of this study is that we have not fully determined in which scenario to use each of our two approaches. Future work aims to expand the number and diversity of tasks to provide better guidance on when to use which method.
In conclusion, we envision that our approaches will create representations that leverage the full range of temporal dynamics and patient history that are typically neglected in current diagnostic and prognostic solutions.